# Uncertainty-Aware Surrogate-based Amortized Bayesian Inference for Computationally Expensive Models

### v2

@@ 1 Introduction
Mathematical models are essential for simulating real-world processes, typically mapping parameters to observable data. Estimating these parameters from real-world observations, i.e., the inverse problem, is fundamental across scientific disciplines (e.g. Etz & Vandekerckhove, 2018; Von Toussaint, 2011; Hülsenbeck et al., 2001; Ellison, 2004). However, because models never perfectly capture reality and observed data are often sparse and imprecise, parameter estimation inherently involves uncertainty. Bayesian inference offers a systematic framework for estimating parameters while incorporating uncertainty in a statistically grounded manner (e.g. Gelman et al., 1995).
@@
Both MCMC and ABI face limitations when the simulation model is computationally expensive, since both require a large number of model evaluations. MCMC requires many likelihood evaluations per generated posterior sample; ABI requires extensive model simulations to generate sufficient amounts of training data. As a result, if computational time is limited, obtaining a good posterior becomes infeasible for both approaches. Note that such computational effort can arise for different reasons – for instance, from a high-dimensional parameter space or from the complexity of the governing equations. In Earth sciences, for example, low-dimensional models can be computationally very expensive, particularly if they involve (coupled systems of) non-linear partial differential equations or long simulated time periods (e.g. Mohammadi et al., 2018; Hommel et al., 2015).
Our goal is to enable ABI for computationally expensive models to benefit from the aforementioned advantages that ABI offers. To this end, surrogate models are the crucial tool, allowing us to reduce the computational cost of expensive simulations for training data generation. Surrogate models (Sudret, 2008; Rasmussen & Williams, 2006), however, are only approximations (i.e., imperfect representations) of the reference simulation, and therefore not necessarily reliable. It is crucial to quantify the uncertainties associated with surrogate modeling, such as those arising from limited training data or any inherent inflexibility of the surrogate. These uncertainties then need to be propagated through the inference pipeline, ensuring consistent posterior estimation. Previous methods have integrated these uncertainties into MCMC methods with surrogate models (e.g. Lückmann et al., 2018; Zhang et al., 2020; Reiser et al., 2025). However, this requires many additional MCMC runs to incorporate the surrogate uncertainty, largely eliminating the computational advantages of surrogates.
We introduce Uncertainty-Aware Surrogate-based Amortized Bayesian Inference (UA-SABI), which enables ABI for computationally expensive models while accounting for the uncertainty inherent in surrogate approximations. By leveraging a surrogate, we eliminate the training error that arises from insufficient data during ABI training, as simulating from the surrogate is easy and fast. This comes at the cost of introducing a surrogate approximation error. However, unlike the ABI training error, the surrogate approximation error can be quantified and propagated, enabling reliable ABI for computationally expensive models.
@@ 2.1 Amortized Bayesian Inference (ABI)
This work focuses on using ABI to infer the parameters of computationally expensive simulation models. The typical ABI procedure (Radev et al., 2020) can be divided into a training phase and an inference phase. Training includes simulating data and then training a neural posterior estimator (NPE) on that data (Papamakarios & Murray, 2016; Durkan et al., 2019; Greenberg et al., 2019b).
@@
After creating the training dataset, the NPE, including a summary network $S_{\bm{\theta}}(\mathbf{x},\mathbf{y})$ and an inference network $I_{\bm{\varphi}}(\mathbf{s})$, must be trained. The goal of $S_{\bm{\theta}}(\mathbf{x},\mathbf{y})$ is to embed (multiple) model inputs and outputs $(\mathbf{x},\mathbf{y})$ of potentially variable length (e.g., a measurement series) in a fixed-length vector $\mathbf{s}$ that serves as input to $I_{\bm{\varphi}}(\mathbf{s})$. Then, $I_{\bm{\varphi}}(\mathbf{s})$ generates samples from the approximate posterior $q_{\bm{\varphi}}(\bm{\omega}\mid\mathbf{s})$. Here, $\bm{\theta}$ and $\bm{\varphi}$ are learnable coefficients in $S_{\bm{\theta}}(\mathbf{x},\mathbf{y})$ and $I_{\bm{\varphi}}(\mathbf{s})$. In our case studies, we use coupling flows (Papamakarios et al., 2021) as inference networks. Coupling flows have been empirically and theoretically shown to be expressive and allow for fast evaluation and sampling (Draxler et al., 2025). However, ABI is generally flexible in its choice for the inference network and works with any generative neural network (e.g. Ardizzone et al., 2019; Padmanabha & Zabaras, 2021; Denker et al., 2021; Wildberger et al., 2023; Schmitt et al., 2024b). Thus, the summary and inference networks are modular and can be independently replaced or adapted to suit the specific problem at hand.
The summary and inference networks are trained jointly by minimizing the Kullback-Leibler (KL) divergence between the true posterior $p(\bm{\omega}\mid\mathbf{x},\mathbf{y})$ and the approximate posterior $q_{\bm{\varphi}}(\bm{\omega}\mid\mathbf{s})$:
$L(\bm{\varphi},\bm{\theta})$ $=\mathbb{E}_{p(\mathbf{y})}[\text{KL}(p(\bm{\omega}\mid\mathbf{x},\mathbf{y})\mid q_{\bm{\varphi}}(\bm{\omega}\mid\mathbf{s}))]\approx\sum_{i=1}^{N_{B}}(-\text{log}~q_{\bm{\varphi}}(\bm{\omega}^{(i)}\mid S_{\bm{\theta}}(\mathbf{x}^{(i)},\mathbf{y}^{(i)})))$ (2)
This expectation is approximated via the training data $D_{B}$ obtained from the (potentially expensive) simulation model (Radev et al., 2020).
@@ 2.2 Uncertainty-Aware Surrogate Modeling
However, due to limited simulation data from $M$ used for training $\widetilde{M}_{\mathbf{c}}$, the surrogate coefficients $\mathbf{c}$ can only be estimated with substantial epistemic uncertainty. For the same reason, the expressibility of $\widetilde{M}_{\mathbf{c}}$ must be restricted, i.e., the complexity of $\widetilde{M}_{\mathbf{c}}$ must be kept relatively low. This introduces an approximation error, denoted as $\epsilon$, which captures the misspecification of the surrogate. In practice, one usually has to expect a non-negligible approximation error due to limitations in model capacity, optimization challenges, or regularization constraints. We assume the approximation error $\epsilon\sim p(\epsilon\mid\sigma)$ to follow a distribution parametrized by $\sigma$, leading to the error-adjusted surrogate output:
$\widetilde{\mathbf{y}}_{\epsilon}$ $=f(\widetilde{\mathbf{y}},\epsilon)\quad\text{with}\quad\epsilon\sim p(\epsilon\mid\sigma).$ (4)
This setup defines the surrogate likelihood
@@
Using sparse training data ($N_{T}\ll N_{B}$) obtained from the simulation model $D_{T}=\{(\mathbf{x}^{(i)},\bm{\omega}^{(i)},\mathbf{y}^{(i)})\}_{i=1}^{N_{T}}$, with $\mathbf{y}^{(i)}=M(\mathbf{x}^{(i)},\bm{\omega}^{(i)})$ and $(\mathbf{x}^{(i)},\bm{\omega}^{(i)})$ sampled from a given prior, we infer the surrogate model coefficients $\mathbf{c}$ and $\sigma$ via Bayesian inference. This enables capturing both the surrogate model’s epistemic uncertainty and the irreducible approximation error. For this purpose, we consider the likelihood equation 5 and the prior $p(\mathbf{c},\sigma)$. The joint posterior over the surrogate parameters and the approximation noise, given the training data, is then
@@
assuming an i.i.d. sampled approximation error. This assumption serves as a simple stand-in, since surrogate construction is not the focus of this work. It suffices to demonstrate that even a basic error model outperforms making no error assumption at all. In general, UA-SABI is not limited to the i.i.d. case; case-specific surrogate error models are fully compatible with the method.
@@ 2.3 Surrogate-based ABI (SABI)
For the point surrogate, the median of the posterior in Eq.˜6 can be used as it is robust to outliers and invariant under monotonic transformations, but also the mean is a valid option.
@@ 2.4 Uncertainty-Aware Surrogate-based ABI (UA-SABI)
To propagate uncertainty to the ABI training, we now additionally draw the surrogate coefficients $\mathbf{c}^{(i)}$ and the approximation error parameter $\sigma^{(i)}$ from the surrogate posterior equation 6. These coefficients are then used to evaluate the surrogate model and obtain $\widetilde{\mathbf{y}}^{(i)}$. The pair $(\widetilde{\mathbf{y}}^{(i)},\sigma^{(i)})$ is subsequently used to generate a sample $\widetilde{\mathbf{y}}_{\epsilon}^{(i)}$ from the distribution of the error-adjusted surrogate output. We now approximate the KL divergence in Eq.˜2 as:
$L(\bm{\varphi},\bm{\theta})\approx\sum_{i=1}^{N_{B}}(-\text{log}~q_{\bm{\varphi}}(\bm{\omega}^{(i)}\mid S_{\bm{\theta}}(\mathbf{x}^{(i)},\widetilde{\mathbf{y}}_{\epsilon}^{(i)})))$ (9)
A pseudocode summarizing the UA-SABI training procedure can be found in Appendix˜A.
@@ 3 Related Work
Multi-fidelity methods leverage models or simulations of varying accuracy and cost to accelerate inference, combining cheap, approximate simulations with a smaller number of expensive, high-fidelity runs. In Approximate Bayesian Computation (ABC), such strategies reduce computational burden: Warne et al. (2018) apply multi-level Monte Carlo variance reduction to rejection sampling in ABC, improving efficiency. Prescott & Baker (2020) introduce a multi-fidelity ABC approach that combines early acceptance and rejection of parameter samples and selectively runs high-fidelity simulations based on low-fidelity outputs, balancing computational cost and accuracy. Building on this, Prescott & Baker (2021) integrate the multi-fidelity strategy with sequential Monte Carlo sampling, further enhancing efficiency by more effectively exploring the parameter space for MF-ABC proposals. More recent work by Warne et al. (2022) integrates low- and high-fidelity simulations within a multilevel Monte Carlo framework to accelerate SBI for partially observed stochastic processes, using inexpensive approximations to preselect parameters and high-fidelity simulations to refine posterior estimates, while Prescott et al. (2024) dynamically select between low- and high-fidelity models based on predictive accuracy, optimizing resource allocation. In multi-fidelity simulation-based Inference (MF-SBI) (Krouglova et al., 2025) a neural posterior estimator is first pre-trained on large amounts of low-fidelity data and then fine-tuned with a smaller set of high-fidelity simulations. While sharing similar goals, none of the aforementioned methods explicitly take the uncertainty of the low-fidelity simulations into account. UA-SABI’s uncertainty-awareness guards against overconfidence in regions with scarce data – a protection which appears not to be in place for the low-fidelity training regions in MF-SBI, for example. Further, unlike many of the multi-fidelity methods, UA-SABI does not require domain expertise for the construction of the surrogate.
@@
S-NPE iteratively refines posterior estimates by adaptively choosing simulations that are most informative, improving efficiency. Papamakarios & Murray (2016) use initial simulations to fit a Bayesian conditional density estimator for the posterior and then sample parameters from the current posterior estimate to guide the next round of simulations. Similarly, Toni et al. (2009) introduced Sequential ABC (ABC-SMC), which iteratively refines posterior distributions by progressively tightening acceptance thresholds. Kulkarni & Moritz (2023) build on ABC-SMC, leveraging massively parallel architectures, such as GPUs, to more efficiently explore the parameter space. Lueckmann et al. (2017) sequentially train a Bayesian mixture-density network. Greenberg et al. (2019a) developed Automatic Posterior Transformation, which employs flexible flow-based density estimators and dynamically updated proposal distributions. Deistler et al. (2022) introduce truncated S-NPE, restricting simulations to truncated regions around high-probability areas. However, UA-SABI and S-NPE pursue related but distinct goals: UA-SABI focuses on fully amortized inference, enabling fast posterior estimation for new datasets through one-time training, whereas S-NPE sequentially refines the posterior on the same dataset to improve accuracy and sample efficiency across rounds. Additionally, the referenced S-NPE procedures implement an active learning strategy for simulation selection, assuming data are generated iteratively, while UA-SABI can also operate on pre-existing simulation datasets (offline training).
@@
Several methods employ Gaussian Process (GP) surrogates to improve sampling efficiency in likelihood-free inference. Meeds & Welling (2014) construct a GP to emulate the expensive simulator and use its predictive mean and uncertainty to estimate the acceptance probability in ABC, avoiding simulator calls when the surrogate is sufficiently confident. Wilkinson (2014) fit a GP to the ABC log-likelihood to prune implausible parameter regions before sampling, greatly reducing simulation cost. Bayesian Optimization likelihood-free inference (Gutmann & Corander, 2016) fits a GP to the discrepancy between simulated and observed data and uses Bayesian optimization to adaptively choose simulator parameters, concentrating effort on informative regions and enabling accurate posterior estimation with few simulations. As in UA-SABI, these approaches apply surrogates, focusing specifically on GPs. However, they employ the surrogate to intelligently select simulation parameters for subsequent likelihood-free inference, whereas UA-SABI uses the surrogate itself to directly generate training data and is agnostic to the surrogate model class.
@@
$p(\bm{\omega}\mid\mathbf{y})=\iint p(\bm{\omega}\mid\mathbf{y},\mathbf{c},\sigma)\,p(\mathbf{c},\sigma\mid D_{T})\,d\mathbf{c}\,d\sigma.$ (10)

### v3

@@ 1 Introduction
Mathematical models are essential for simulating real-world processes, typically mapping parameters to observable data. Estimating these parameters from real-world observations, i.e., the inverse problem, is fundamental across scientific disciplines (e.g. Etz and Vandekerckhove, 2018; Von Toussaint, 2011; Hülsenbeck et al., 2001; Ellison, 2004). However, because models never perfectly capture reality and observed data are often sparse and imprecise, parameter estimation inherently involves uncertainty. Bayesian inference offers a systematic framework for estimating parameters while incorporating uncertainty in a statistically grounded manner (e.g. Gelman et al., 1995).
@@
Both MCMC and ABI face limitations when the simulation model is computationally expensive, since both require a large number of model evaluations. MCMC requires many likelihood evaluations per generated posterior sample; ABI requires extensive model simulations to generate sufficient amounts of training data. As a result, if computational time is limited, obtaining a good posterior becomes infeasible for both approaches. Note that such computational effort can arise for different reasons – for instance, from a high-dimensional parameter space or from the complexity of the governing equations. In Earth sciences, for example, low-dimensional models can be computationally very expensive, particularly if they involve (coupled systems of) non-linear partial differential equations or long simulated time periods (e.g. Mohammadi et al., 2018; Hommel et al., 2015), resulting in runtimes up to multiple days for a single model run.
Our goal is to enable ABI for computationally expensive models to benefit from the aforementioned advantages that ABI offers. To this end, surrogate models are the crucial tool, allowing us to reduce the computational cost of expensive simulations for training data generation. Surrogate models (Sudret, 2008; Rasmussen and Williams, 2006), however, are only approximations (i.e., imperfect representations) of the reference simulation, and therefore not necessarily reliable. It is crucial to quantify the uncertainties associated with surrogate modeling, such as those arising from limited training data or any inherent inflexibility of the surrogate. These uncertainties then need to be propagated through the inference pipeline, ensuring consistent posterior estimation. Previous methods have integrated these uncertainties into MCMC methods with surrogate models (e.g. Lückmann et al., 2018; Zhang et al., 2020; Reiser et al., 2025). However, this requires many additional MCMC runs to incorporate the surrogate uncertainty, largely eliminating the computational advantages of surrogates.
We introduce Uncertainty-Aware Surrogate-based Amortized Bayesian Inference (UA-SABI), which enables ABI for computationally expensive models while accounting for the uncertainty inherent in surrogate approximations. By leveraging a surrogate, we reduce the training error that arises from insufficient data during ABI training, as simulating from the surrogate is easy and fast. This comes at the cost of introducing a surrogate approximation error. However, unlike the ABI training error, the surrogate approximation error can be quantified and propagated, enabling reliable ABI for computationally expensive models.
@@ 2.1 Amortized Bayesian Inference (ABI)
This work focuses on using ABI to infer the parameters of computationally expensive simulation models. The typical ABI procedure (Radev et al., 2020) can be divided into a training phase and an inference phase. Training includes simulating data and then training a neural posterior estimator (NPE) on that data (Papamakarios and Murray, 2016; Durkan et al., 2019; Greenberg et al., 2019b).
@@
After creating the training dataset, the NPE, including a summary network $S_{\bm{\theta}}(\mathbf{x},\mathbf{y})$ and an inference network $I_{\bm{\varphi}}(\mathbf{s})$, must be trained. The goal of $S_{\bm{\theta}}(\mathbf{x},\mathbf{y})$ is to embed (stacked) observations, i.e., model inputs and outputs $(\mathbf{x},\mathbf{y})$ of potentially variable length (e.g., a measurement series) in a fixed-length vector $\mathbf{s}$ that serves as input to $I_{\bm{\varphi}}(\mathbf{s})$. Then, $I_{\bm{\varphi}}(\mathbf{s})$ generates samples from the approximate posterior $q_{\bm{\varphi}}(\bm{\omega}\mid\mathbf{s})$. Here, $\bm{\theta}$ and $\bm{\varphi}$ are learnable coefficients in $S_{\bm{\theta}}(\mathbf{x},\mathbf{y})$ and $I_{\bm{\varphi}}(\mathbf{s})$. In our case studies, we use coupling flows (Papamakarios et al., 2021) as inference networks. Coupling flows have been empirically and theoretically shown to be expressive and allow for fast evaluation and sampling (Draxler et al., 2025). However, ABI is generally flexible in its choice for the inference network and works with any generative neural network (e.g. Ardizzone et al., 2019; Padmanabha and Zabaras, 2021; Denker et al., 2021; Wildberger et al., 2023; Schmitt et al., 2024b). Thus, the summary and inference networks are modular and can be independently replaced or adapted to suit the specific problem at hand.
The optimal parameters of the summary and inference network $(\bm{\theta}^{*},\bm{\varphi}^{*})$ are obtained jointly by minimizing the Kullback-Leibler (KL) divergence between the true posterior $p(\bm{\omega}\mid\mathbf{x},\mathbf{y})$ and the approximate posterior $q_{\bm{\varphi}}(\bm{\omega}\mid\mathbf{s})$ (Radev et al., 2020):
$(\bm{\theta}^{*},\bm{\varphi}^{*})$ $=\underset{\bm{\theta},\bm{\varphi}}{\text{argmin}}\,\mathbb{E}_{p(\mathbf{x},\mathbf{y})}[\text{KL}(p(\bm{\omega}\mid\mathbf{x},\mathbf{y})\mid\mid q_{\bm{\varphi}}(\bm{\omega}\mid\mathbf{s}))]$ (2)
$=\underset{\bm{\theta},\bm{\varphi}}{\text{argmin}}\,\mathbb{E}_{p(\mathbf{x},\mathbf{y})}[\mathbb{E}_{p(\bm{\omega}\mid\mathbf{x},\mathbf{y})}[-\text{log}\,q_{\bm{\varphi}}(\bm{\omega}\mid\mathbf{s})]]$ (3)
$=\underset{\bm{\theta},\bm{\varphi}}{\text{argmin}}\,-\iiint p(\mathbf{x},\mathbf{y},\bm{\omega})\,\text{log}\,q_{\bm{\varphi}}(\bm{\omega}\mid S_{\bm{\theta}}(\mathbf{x},\mathbf{y}))\,\text{d}\mathbf{x}\,\text{d}\mathbf{y}\,\text{d}\bm{\omega}.$ (4)
According to Radev et al. (2020), the expectations can be approximated with the Monte Carlo estimates, utilizing simulations $\{(\mathbf{x}^{(i)},\bm{\omega}^{(i)},\mathbf{y}^{(i)})\}_{i=1}^{N_{B}}$ from the (potentially expensive) simulation model (1). This results in the following loss function:
$L(\bm{\varphi},\bm{\theta})=-\sum_{i=1}^{N_{B}}\text{log}\,q_{\bm{\varphi}}(\bm{\omega}^{(i)}\mid S_{\bm{\theta}}(\mathbf{x}^{(i)},\mathbf{y}^{(i)})).$ (5)
@@ 2.2 Uncertainty-Aware Surrogate Modeling
However, due to limited simulation data from $M$ used for training $\widetilde{M}_{\mathbf{c}}$, the surrogate coefficients $\mathbf{c}$ can only be estimated with substantial epistemic uncertainty. For the same reason, the expressibility of $\widetilde{M}_{\mathbf{c}}$ must be restricted, i.e., the complexity of $\widetilde{M}_{\mathbf{c}}$ must be kept relatively low. This introduces an approximation error, denoted as $\epsilon$, which captures the misspecification of the surrogate. In practice, one usually has to expect a non-negligible approximation error due to limitations in model capacity, optimization challenges, or regularization constraints. We assume the approximation error $\epsilon\sim p(\epsilon\mid\sigma)$ to follow a distribution parametrized by $\sigma$, leading to the error-adjusted surrogate output
$\widetilde{\mathbf{y}}_{\epsilon}$ $=f(\widetilde{\mathbf{y}},\epsilon)\quad\text{with}\quad\epsilon\sim p(\epsilon\mid\sigma),$ (7)
which is essentially a perturbed version of $\widetilde{\mathbf{y}}$. This setup defines the surrogate likelihood
@@
Using sparse training data ($N_{T}\ll N_{B}$) obtained from the simulation model $D_{T}=\{(\mathbf{x}^{(i)},\bm{\omega}^{(i)},\mathbf{y}^{(i)})\}_{i=1}^{N_{T}}$, with $\mathbf{y}^{(i)}=M(\mathbf{x}^{(i)},\bm{\omega}^{(i)})$ and $(\mathbf{x}^{(i)},\bm{\omega}^{(i)})$ sampled from a given prior, we infer the surrogate model coefficients $\mathbf{c}$ and $\sigma$ via Bayesian inference. This enables capturing both the surrogate model’s epistemic uncertainty and the irreducible approximation error. For this purpose, we consider the likelihood equation 8 and the prior $p(\mathbf{c},\sigma)$. The joint posterior over the surrogate parameters and the approximation noise, given training data $D_{T}$ from the simulation model $M$, is then
@@
assuming an i.i.d. sampled approximation error. This assumption serves as a simple stand-in, since surrogate construction is not the focus of this work. It suffices to demonstrate that even a basic error model outperforms making no error assumption at all. In general, UA-SABI is not limited to the i.i.d. case; heteroscedastic surrogate error models (e.g. Kohlhaas et al., 2023) are fully compatible with the method.
@@ 2.3 Surrogate-based ABI (SABI)
For the point surrogate, the median of the posterior in Eq.˜9 can be used as it is robust to outliers and invariant under monotonic transformations, but also the mean is a valid option.
@@ 2.4 Uncertainty-Aware Surrogate-based ABI (UA-SABI)
To propagate uncertainty to the ABI training, we now additionally draw the surrogate coefficients $\mathbf{c}^{(i)}$ and the approximation error parameter $\sigma^{(i)}$ from the surrogate posterior equation 9. These coefficients are then used to evaluate the surrogate model and obtain $\widetilde{\mathbf{y}}^{(i)}$. The pair $(\widetilde{\mathbf{y}}^{(i)},\sigma^{(i)})$ is subsequently used to generate a sample $\widetilde{\mathbf{y}}_{\epsilon}^{(i)}$ from the distribution of the error-adjusted surrogate output. We now approximate the KL divergence in Eq.˜5 as:
$L(\bm{\varphi},\bm{\theta})=-\sum_{i=1}^{N_{B}}\text{log}\,q_{\bm{\varphi}}(\bm{\omega}^{(i)}\mid S_{\bm{\theta}}(\mathbf{x}^{(i)},\widetilde{\mathbf{y}}_{\epsilon}^{(i)})).$ (12)
Figure˜2 gives a detailed graphical overview. Additionally, a pseudocode summarizing the UA-SABI training procedure can be found in Appendix˜A.
Figure 2: Detailed graphical overview of UA-SABI training as implemented in Algorithm˜1. First phase: training of uncertainty-aware surrogate model using sparse simulation data, yielding a posterior over surrogate parameters. Second phase: ABI training using surrogate-generated data, where full surrogate uncertainty is propagated through sampled surrogate outputs.
@@ 3 Related Work
Multi-fidelity methods leverage models or simulations of varying accuracy and cost to accelerate inference, combining cheap, approximate simulations with a smaller number of expensive, high-fidelity runs. In Approximate Bayesian Computation (ABC), such strategies reduce computational burden: Warne et al. (2018) apply multi-level Monte Carlo variance reduction to rejection sampling in ABC, improving efficiency. Prescott and Baker (2020) introduce a multi-fidelity ABC approach that combines early acceptance and rejection of parameter samples and selectively runs high-fidelity simulations based on low-fidelity outputs, balancing computational cost and accuracy. Building on this, Prescott and Baker (2021) integrate the multi-fidelity strategy with sequential Monte Carlo sampling, further enhancing efficiency by more effectively exploring the parameter space for MF-ABC proposals. More recent work by Warne et al. (2022) integrates low- and high-fidelity simulations within a multilevel Monte Carlo framework to accelerate SBI for partially observed stochastic processes, using inexpensive approximations to preselect parameters and high-fidelity simulations to refine posterior estimates, while Prescott et al. (2024) dynamically select between low- and high-fidelity models based on predictive accuracy, optimizing resource allocation. In multi-fidelity simulation-based Inference (MF-SBI) (Krouglova et al., 2025) a neural posterior estimator is first pre-trained on large amounts of low-fidelity data and then fine-tuned with a smaller set of high-fidelity simulations. While sharing similar goals, none of the aforementioned methods explicitly take the uncertainty of the low-fidelity simulations into account. UA-SABI’s uncertainty-awareness guards against overconfidence in regions with scarce data – a protection which appears not to be in place for the low-fidelity training regions in MF-SBI, for example. Further, unlike many of the multi-fidelity methods, UA-SABI does not require domain expertise for the construction of the surrogate.
@@
S-NPE iteratively refines posterior estimates by adaptively choosing simulations that are most informative, improving efficiency. Papamakarios and Murray (2016) use initial simulations to fit a Bayesian conditional density estimator for the posterior and then sample parameters from the current posterior estimate to guide the next round of simulations. Similarly, Toni et al. (2009) introduced Sequential ABC (ABC-SMC), which iteratively refines posterior distributions by progressively tightening acceptance thresholds. Kulkarni and Moritz (2023) build on ABC-SMC, leveraging massively parallel architectures, such as GPUs, to more efficiently explore the parameter space. Lueckmann et al. (2017) sequentially train a Bayesian mixture-density network. Greenberg et al. (2019a) developed Automatic Posterior Transformation, which employs flexible flow-based density estimators and dynamically updated proposal distributions. Deistler et al. (2022) introduce truncated S-NPE, restricting simulations to truncated regions around high-probability areas. However, UA-SABI and S-NPE pursue related but distinct goals: UA-SABI focuses on fully amortized inference, enabling fast posterior estimation for new datasets through one-time training, whereas S-NPE sequentially refines the posterior on the same dataset to improve accuracy and sample efficiency across rounds. Additionally, the referenced S-NPE procedures implement an active learning strategy for simulation selection, assuming data are generated iteratively, while UA-SABI can also operate on pre-existing simulation datasets (offline training).
@@
Several methods employ Gaussian Process (GP) surrogates to improve sampling efficiency in likelihood-free inference. Meeds and Welling (2014) construct a GP to emulate the expensive simulator and use its predictive mean and uncertainty to estimate the acceptance probability in ABC, avoiding simulator calls when the surrogate is sufficiently confident. Wilkinson (2014) fit a GP to the ABC log-likelihood to prune implausible parameter regions before sampling, greatly reducing simulation cost. Bayesian Optimization likelihood-free inference (Gutmann and Corander, 2016) fits a GP to the discrepancy between simulated and observed data and uses Bayesian optimization to adaptively choose simulator parameters, concentrating effort on informative regions and enabling accurate posterior estimation with few simulations. As in UA-SABI, these approaches apply surrogates, focusing specifically on GPs. However, they employ the surrogate to intelligently select simulation parameters for subsequent likelihood-free inference, whereas UA-SABI uses the surrogate itself to directly generate training data and is agnostic to the surrogate model class.
@@
$p(\bm{\omega}\mid\mathbf{y})=\iint p(\bm{\omega}\mid\mathbf{y},\mathbf{c},\sigma)\,p(\mathbf{c},\sigma\mid D_{T})\,\text{d}\mathbf{c}\,\text{d}\sigma.$ (13)